OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = -31 and a9 = -119? (1 points) -372 -361 -350 -339

OpenStudy (anonymous):

an=a1+(n-1)d => a9=a1+(9-1)d => 8d=a9-a1=-119-(-31)=-88 => d=-11. then a32=a1+(32-1)d=-31+31*-11=-372. therefore a32=-372

OpenStudy (anonymous):

Thank You So Much

OpenStudy (anonymous):

no problem. any way, is it clear?

OpenStudy (anonymous):

Yes Thanks